OpenStudy (anonymous):

Write the formula for the geometric sequence. 5, 10, 20, 40, … A. an=40∙2n-1 B. an=40∙2n C. an=5∙2n-1 D. an=5∙2n

OpenStudy (anonymous):

I have no idea how to do this. It would be great if someone could teach me!

OpenStudy (anonymous):

a=5,r=10/5=2 \[an=a*r ^{n-1}\]

OpenStudy (***[isuru]***):

the general format of a geometric progression is \[T_{n} = ar^{n-1}\] where a = the first term n = nth term r = common ration u already know that the the first term is 5 ...so the a= 5 now to find (r) if there is 2 consecutive terms as b , c \[ r = \frac{ c }{ b }\] now if u take b = 5 c= 10 what is r ?

OpenStudy (anonymous):

So @***[ISURU]*** is r 0.5?